AT&T Inc. had the EBIT (earnings before interest and tax) of $31,618 million, interest expense of $7,959 million, and the tax expense of $4,920 million. Its current share price is $30.99 per share. It has 7,280 million shares outstanding. Compute its current Price-Earnings (P/E) ratio,
Answer #1
EBIT 31618 million
Less:interest 7959 million
EBT 23659 million
Less:tax expense 4920 million
Net income 18739 million

EPS=Net income/Shares outstanding

=(18739/7280)=$2.574038462

PE ratio=Market price/EPS

=(30.99/2.574038462)

=12.04(Approx).